# The spectral genus of an isolated hypersurface singularity and a conjecture relating to the Milnor number

### v1

@@ Abstract.
In this paper, we introduce the notion of spectral genus $\widetilde{p}_{g}$ of a deformation of a germ of an isolated hypersurface singularity $(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, defined as a sum of small exponents of monodromy eigenvalues. The number of these is equal to the geometric genus $p_{g}$, and hence $\widetilde{p}_{g}$ can be considered as a secondary invariant to it. We then explore a secondary version of the Durfee conjecture on $p_{g}$, regarding the relationship between $\widetilde{p}_{g}$ the Milnor number $\mu$, to the effect that
$\frac{\mu}{(n+2)!}>\widetilde{p}_{g}.$
We provide evidence by confirming the conjecture in several cases, including homogeneous singularities and singularities with large Newton polyhedra, and quasi-homogeneous or irreducible curve singularities.
@@
Isolated hypersurface singularities, Spectrum of singularities, Milnor numbers
@@ 1. Introduction
In 1978, Durfee conjectured an inequality between the geometric genus $p_{g}$ of complete intersection isolated surface singularities and their Milnor numbers $\mu$, in [11]. While the initial expectation does not hold in such generality, the conjecture was later generalized in [27] by K. Saito to a conjecture about general germs of $n$-dimensional isolated hypersurface singularities [^1] In this paper, by a singularity, we mean a non-regular point., defined by germs of holomorphic functions $f:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, as an inequality
(1.1) $\frac{\mu}{(n+1)!}>p_{g}.$
This generalized inequality appeared naturally in his investigation of the distribution of the spectrum of the semi-simple part of the monodromy in the same article. Particular cases and variants of the Durfee-type conjecture (1.1) have been established by Yau–Zhang [36] and Kerner–Némethi [17, 18].
@@
where the sum is over the rational numbers $\lambda_{j}\in[0,1)$ such that $\exp(2\pi i\lambda_{j})$ is an eigenvalue of the semi-simple part of the monodromy acting on $\operatorname{Gr}^{n}_{F}H^{n}(\operatorname{Mil}_{f})$. It is a sort of secondary invariant of the geometric genus, and is in particular zero for rational singularities. One may wonder if there is a corresponding secondary version of Durfee’s conjecture. In light of this, we propose the following:
@@
For an $n$-dimensional isolated hypersurface singularity, we have an inequality:
@@
    $\frac{\mu}{(n+2)!}\geq\widetilde{p}_{g}+\frac{1}{(n+2)!}.$
The inequality in the strong form is attained for ordinary double points, since in this case the spectral genus is zero and the Milnor number is one. It seems natural to wonder if this is the only case where the inequality is attained. Moreover, if K. Saito’s conjectures in [27] about the distribution of monodromy eigenvalues hold, then the above inequalities are essentially optimal. Namely, it implies (cf. Proposition 7.2) that
@@
Our contribution is the following partial confirmation:
Main Theorem.
@@
The extension of the conjecture to higher dimensions was partially motivated by the asymptotic behavior of the holomorphic analytic torsion, which is defined in terms of regularized determinants of Dolbeault–Laplace operators. More precisely, the behavior of $\tau(\mathcal{O}_{\mathcal{X}_{t}})$ for a degenerating family of projective varieties $\mathcal{X}\to\Delta$ with isolated singularities, is captured by the expression $(-1)^{n}(\mu/(n+2)!-\widetilde{p}_{g})$ (cf. Corollary 3.2). This is related to the work of Yoshikawa on the singularities of the Quillen metric [37, 38]. See also [12].
@@
is continuous over the space of hypersurfaces in $\mathbb{P}^{n+1}$. More precisely, the weak form of the conjecture indicates that the analytic torsion even vanishes along the discriminant locus. In fact, this continuity statement implies the weak form of the conjecture (cf. §2.7 and §3.7). This is based on the fact the function in (1.2), if continuous, vanishes on the locus of ordinary double points because the weak form of the conjecture is known then, and any isolated singularity is a limit of ordinary double points. This formulation of the problem possibly invites analytic techniques, of global nature, to study the conjecture, which is of local type.
@@ 2. Invariants of isolated singularties
In this section, to set up notation and for the convenience of the reader, we recall some classical invariants of isolated hypersurface singularities. We will be considering the germ of a holomorphic function $f:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$ defining an isolated hypersurface singularity at the origin, and we will suppose that $n\geq 1$. Somewhat abusively, in this text we will simply write $f=0$.
@@
$\mu=\dim\mathbb{C}\{z_{0},\ldots,z_{n}\}/\left(\partial f/\partial z_{0},\ldots,\partial f/\partial z_{n}\right).$
The cohomology group $H^{n}(\operatorname{Mil}_{f})$ is equipped with a canonical mixed Hodge structure by [32]. We denote the corresponding Hodge filtration by $F$. Moreover, the semi-simple part of the monodromy, denoted by $T_{s}$, acts on this mixed Hodge, in the sense that it preserves the weight and Hodge filtrations [32, Theorem 4.1].
@@
Consider the the genus of the singularity $V=\{f=0\}$:
@@
where $\pi:X\to V$ is a desingularization. The equality of the various quantities is proven in [28, Theorem 1] and [33, Proposition 2.13]. Recall from the introduction that we likewise define the spectral genus of the singularity, as the expression
@@
We describe the relationship between the spectral genus and the spectrum of an isolated hypersurface singularity. We first follow Steenbrink’s presentation in [34, Section 2] and [35, Section 1], and we adopt his conventions. See also [25, Section 12.1.3]. The spectral numbers are associated to the triple $(H^{n}(\operatorname{Mil}_{f}),F^{\bullet},T_{s})$. These are rational numbers $\alpha$, given with multiplicities, uniquely determined by the following conditions:
@@
$\widetilde{p}_{g}=-\sum_{-1<\alpha_{j}<0}\alpha_{j}.$
@@
(2.4) $\widetilde{p}_{g}=\sum_{0<\alpha_{j}^{\prime}<1}(1-\alpha_{j}^{\prime}).$
For instance, this is the case of the works of M. Saito [29, 30] used below. Depending on the context, one convention may be more adapted than the other, and we will use both.
@@
2.4.
The Milnor number $\mu$ and the geometric genus $p_{g}$ depend only on the fiber $V=f^{-1}(0)$, but not on the chosen deformation [34, Theorem 2.9]. The spectral genus depends on the whole germ $f$. Nevertheless, the spectrum is constant in any deformation of isolated hypersurface singularities with constant $\mu$ [34, Theorem 2.8], and hence so is the spectral genus.
2.5.
Consider a connected complex manifold $\mathcal{X}$ of dimension $n+1$, and a flat, projective morphism $g:\mathcal{X}\to\Delta$, which is a holomorphic submersion outside the origin. We will suppose that the central fiber has at most isolated singularities. If $x_{i}\in\mathcal{X}_{0}$ is such a singular point, the germ $(\mathcal{X},x_{i})\to(\Delta,0)$ is isomorphic to some $f_{i}:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$ and admits a Milnor fiber $\operatorname{Mil}_{f_{i}}$. In this setting, we denote by $\mu$ and $\widetilde{p}_{g}$ the sum of the Milnor numbers and spectral genera of the singularities $x_{i}$, respectively. Therefore, there is a decomposition
@@
and for $q\neq n,n+1$ there is an isomorphism
(2.7) $H^{q}(\mathcal{X}_{0})\simeq H^{q}(\mathcal{X}_{t}),$
where $t\in\Delta\setminus\{0\}$. When $H^{n}(\mathcal{X}_{0})$ is given the canonical mixed Hodge structure of Deligne, and $H^{n}(\mathcal{X}_{t})$ is given the limit mixed Hodge structure, this is moreover an exact sequence of mixed Hodge structures. It is equivariant with respect to the semi-simple part of the monodromy, $T_{s}$. For details, we refer to Steenbrink [32, Section 3.3] and Navarro Aznar [23, Section 14].
2.7.
@@
By an argument of Brieskorn [6, Section 1.1], such a compactification exists, and one can further suppose that it is given by a family of hypersurfaces in $\mathbb{P}^{n+1}$. This moreover shows that properties such as the positivity of $\frac{\mu}{(n+2)!}-\widetilde{p}_{g}$ can be studied equivalently for proper families or in the local setting.
@@ 3. Analytic torsion of $\mathcal{O}_{X}$
In this section, we discuss the asymptotic behavior of the holomorphic analytic torsion of the sheaf of holomorphic functions, for a degeneration of projective varieties.
@@
More generally, for a flat proper morphism of complex analytic spaces $\mathcal{X}\to S$ over a complex analytic manifold $S$, and a vector bundle $E$ on $\mathcal{X}$, there is a line bundle $\lambda(E)$ on $S$, whose fibers over $s\in S$ are given by $\lambda(E|_{\mathcal{X}_{s}})$, see [3, Section 4.1]. It is also referred to as the determinant of the cohomology. If $\mathcal{X}\to S$ is the analytification of an algebraic family, this construction is the analytification of the Knudsen–Mumford determinant [19]. It has a natural grading, which for the purposes of this article, together with various sign issues, we can ignore.
@@
Let $X$ be a compact Kähler manifold. The analytic torsion of a hermitian vector bundle $E$ on $X$ is given by
$\tau(X,E)=\exp\left(\sum(-1)^{q+1}q\zeta^{\prime}_{0,q}(0)\right)$
where, for $\operatorname{Re}(s)\gg 0$, $\zeta_{0,q}(s)$ is given by
@@
whose sum runs over positive eigenvalues $\mu_{j}$ of the Dolbeault-Laplacian $\Delta_{\overline{\partial}}$ acting on $A^{0,q}(E)$. It depends on both the Kähler metric on $X$ and the hermitian metric on $E$.
@@
Let $\mathcal{X}\to\Delta$ be a degeneration of projective varieties of dimension $n$ as in §2.7 (1), with at most isolated singularities in the central fiber $\mathcal{X}_{0}$. Suppose that $\mathcal{X}$ is equipped with a hermitian metric, which is fiberwise Kähler, and that we are given a hermitian vector bundle $E$ on $\mathcal{X},$ of rank $e$. Let $\sigma$ be a holomorphic trivialization of the determinant of the cohomology $\lambda(E)$. By the main result of [37], we have
(3.3) $\log\|\sigma\|_{Q}^{2}=\frac{(-1)^{n}}{(n+2)!}\mu\cdot e\cdot\log|t|^{2}+o(\log|t|),\quad\text{as}\quad t\to 0.$
@@
Suppose now that $\mathcal{X}\to\Delta$ is as in §3.4, but $\mathcal{X}$ is actually equipped with a Kähler metric, whose associated Kähler form is rational when restricted to smooth fibers.
@@
$\log\|\sigma\|^{2}_{L^{2}}=(-1)^{n}\widetilde{p}_{g}\log|t|^{2}+o(\log(t)).$
@@
First of all, the determinant of the cohomology of the sheaf of holomorphic functions is invariant under blowups in the special fiber, because the total space is smooth and hence has only rational singularities. We can hence suppose the central fiber of $\mathcal{X}\to\Delta$ has normal crossings. In this case, by [13, Theorem C] we find that
$\log\|\sigma\|_{L^{2}}^{2}=\left(\sum(-1)^{q}\alpha^{0,q}\right)\log|t|^{2}+o(\log|t|),$
where $\alpha^{0,q}$ is defined as minus (the lower extension of) the logarithm of the semi-simple part of the monodromy acting on $H^{0,q}(\mathcal{X}_{t})$. One can infer directly from the definition that $\alpha^{0,n}=\widetilde{p}_{g}$. Moreover, $\alpha^{0,q}=0$ otherwise. This latter fact follows from (2.6) and (2.7), paired with the fact that $H^{q}(\mathcal{X}_{0})$ has trivial monodromy for any $q$. ∎
@@
Let $\mathcal{X}\to\Delta$ be a degeneration as in §3.5. We conclude by §3.3 and Proposition 3.1 the following:

### v3

@@ Abstract.
In this paper, we introduce the notion of spectral genus $\widetilde{p}_{g}$ of a germ of an isolated hypersurface singularity $(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, defined as a sum of small exponents of monodromy eigenvalues. The number of these is equal to the geometric genus $p_{g}$, and hence $\widetilde{p}_{g}$ can be considered as a secondary invariant to it. We then explore a secondary version of the Durfee conjecture on $p_{g}$, and we predict an inequality between $\widetilde{p}_{g}$ and the Milnor number $\mu$, to the effect that
$\widetilde{p}_{g}\leq\frac{\mu-1}{(n+2)!}.$
We provide evidence by confirming our conjecture in several cases, including homogeneous singularities and singularities with large Newton polyhedra, and quasi-homogeneous or irreducible curve singularities. We also show that a weaker inequality follows from Durfee’s conjecture, and hence holds for quasi-homogeneous singularities and curve singularities.
@@
Isolated hypersurface singularities, Spectrum of singularities, Milnor numbers, Analytic torsion
@@ 1. Introduction
In 1978, Durfee conjectured an inequality between the geometric genus $p_{g}$ and the Milnor number of complete intersection isolated surface singularities, in [13]. While the initial expectation does not hold in such generality, the conjecture was later extended in [32] by K. Saito to a conjecture about general germs of $n$-dimensional isolated hypersurface singularities [^1] In this paper, by a singularity, we mean a non-regular point., with $n\geq 2$, defined by germs of holomorphic functions $f:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, as an inequality
(1.1) $p_{g}<\frac{\mu}{(n+1)!}.$
This generalized inequality appeared naturally in his investigation of the distribution of the spectrum of the semi-simple part of the monodromy, in the same article. Particular cases and variants of the Durfee-type conjecture (1.1) have since been established by Némethi [27, 29] (suspension-type surfaces), Yau–Zhang [45] (quasi-homogeneous singularities) and Kerner–Némethi [21] (generic singularities with large Newton diagram).
@@
where the sum is over the rational numbers $\lambda_{j}\in[0,1)$ such that $\exp(2\pi i\lambda_{j})$ is an eigenvalue of the semi-simple part of the monodromy acting on $\operatorname{Gr}^{n}_{F}H^{n}(\operatorname{Mil}_{f})$. It is a sort of secondary invariant of the geometric genus, and is in particular zero for rational or, equivalently, canonical singularities. One may wonder if there is a corresponding secondary version of the Durfee-type conjecture. In light of this, we propose the following:
@@
For an $n$-dimensional isolated hypersurface singularity, with $n\geq 1$, we have an inequality:
@@
    $\widetilde{p}_{g}\leq\frac{\mu-1}{(n+2)!}.$
The inequality in the strong form is clearly satisfied for rational singularities, since then the spectral genus vanishes and $\mu\geq 1$. It is attained in the particular case of ordinary double points, since the Milnor number is moreover one. It seems natural to wonder if this is the only situation where the inequality is attained. Moreover, if K. Saito’s conjectures in [32] about the distribution of monodromy eigenvalues hold, then the above inequalities are essentially optimal. Namely, it implies (cf. Proposition 7.4) that
@@
Our first main contribution is the following partial confirmation:
Theorem A.
@@
The extension of the conjecture to higher dimensions was motivated by the asymptotic behavior of the holomorphic analytic torsion, which is defined in terms of regularized determinants of Dolbeault–Laplace operators. It is denoted by $\tau$, and it is a strictly positive real number. More precisely, the rate of vanishing or blowing-up of $\tau(\mathcal{O}_{\mathcal{X}_{t}})$ for a degenerating family of projective varieties $\mathcal{X}\to\Delta$ with isolated singularities, is essentially captured by the expression $(-1)^{n}(\mu/(n+2)!-\widetilde{p}_{g})$ (cf. Corollary 3.2). This is related to the work of Yoshikawa on the singularities of the Quillen metric [46, 47]. See also [15] for an interpretation of Yoshikawa’s results in terms of intersection theory.
@@
extends continuously over the space of hypersurfaces in $\mathbb{P}^{n+1}$. Such a property would imply the weak form of the conjecture (cf. §3.6 to §3.8). This is based on the fact that the function in (1.2), if continuous, vanishes on the locus of ordinary double points, because the weak form of the conjecture is known then, and any isolated singularity is a limit of ordinary double points. The strong form of the conjecture even allows to predict the order of vanishing along the discriminant locus. This formulation of the problem possibly invites analytic techniques of global nature to study the conjecture, which is of local type.
@@
By means of a suspension trick, that J. I. Burgos Gil generously shared with us, we prove in Proposition 7.2 that the weak form of the conjecture is a consequence of the Durfee-type conjecture (1.1). This leads to the second main result of this article:
Theorem B.
The weak form of the conjecture holds in the following cases:
  1. (1)
    Plane curves singularities, ie. when $n=1$ in general.
  2. (2)
    Quasi-homogeneous singularities in arbitrary dimension.
The proof, presented in §7.2, is deduced from the work of Némethi on the Durfee conjecture for suspensions of curves, and the work of Yau–Zhang for quasi-homogeneous singularities, both recalled in §1.1 above.
As for the strong form of the conjecture, it would be interesting to relate it to a Durfee-type bound as well, but we were not able to derive it from the existing refinements, such as those in [21, 44, 45].
1.6.
Exploiting the relationship between our conjecture and the asymptotic behavior of the analytic torsion stated in Corollary 3.2, Theorem A and Theorem B have the following concrete application to determinants of Laplacians, which appears to be new.
Theorem C.
Let $\mathcal{X}\to\Delta$ a degeneration of compact Riemann surfaces, with isolated singularities in the central fiber $\mathcal{X}_{0}$ and with $\mathcal{X}$ smooth Kähler. Then, the determinant of the Laplace–Beltrami operator on $\mathcal{X}_{t}$ satisfies
$\det\Delta_{\mathcal{X}_{t}}\to 0,\quad\text{as}\quad t\to 0.$
Furthermore, if the singularities of $\mathcal{X}_{0}$ are locally irreducible or quasi-homogeneous, then for every $\varepsilon>0$ we have
$\det\Delta_{\mathcal{X}_{t}}=O\left(|t|^{m/3-\varepsilon}\right),$
where $m$ is the number of singular points in $\mathcal{X}_{0}$.
The strong form of the conjecture suggests that the second part of Theorem C should hold for general projective degenerations with isolated singularities.
For degenerations of compact hyperbolic Riemann surfaces, an analogue of Theorem C above is known too. This is contained in the work of Wolpert on degenerations of Selberg zeta functions [43], and can also be derived from [19]. In this setting, it is enough to consider stable degenerations, in which case the singular fiber is endowed with a complete metric. This is somewhat opposite to the situation treated in Theorem C, for which it seems not possible to perform a semi-stable reduction, and moreover the singular fiber carries an incomplete metric.
1.7.
@@ 2. Invariants of isolated singularties
In this section, to set up notation and for the convenience of the reader, we recall some classical invariants of isolated hypersurface singularities. We will be considering the germ of a holomorphic function $f:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$ defining an isolated hypersurface singularity at the origin, and we will suppose that $n\geq 1$. Somewhat abusively, sometimes we will simply write $f=0$.
@@
$\mu=\mu_{f}=\dim\mathbb{C}\{x_{0},\ldots,x_{n}\}/\left(\partial f/\partial x_{0},\ldots,\partial f/\partial x_{n}\right).$
The cohomology group $H^{n}(\operatorname{Mil}_{f})$ is equipped with a canonical mixed Hodge structure by [37]. We denote the corresponding Hodge and weight filtrations by $F$ and $W$, respectively. Moreover, the semi-simple part of the monodromy, denoted by $T_{s}$, acts on this mixed Hodge, in the sense that it preserves the weight and Hodge filtrations [37, Theorem 4.1].
@@
Consider the genus of the singularity $V=\{f=0\}$:
@@
where $\pi:X\to V$ is a desingularization. The equality of the various quantities is proven in [33, Theorem 1] and [38, Proposition 2.13]. In particular, $p_{g}=0$ if $V$ has rational singularities. In our setting, since $V$ is Gorenstein, this is equivalent to saying that $V$ has canonical singularities [14].
Recall from the introduction that we likewise define the spectral genus of the singularity, as the expression
@@
We next elaborate on the relationship between the spectral genus and the spectrum of an isolated hypersurface singularity. We first follow Steenbrink’s presentation in [39, Section 2] and [40, Section 1], and we adopt his conventions. See also [30, Section 12.1.3]. The spectral numbers are associated to the triple $(H^{n}(\operatorname{Mil}_{f}),F^{\bullet},T_{s})$. These are rational numbers $\alpha$, given with multiplicities, uniquely determined by the following conditions:
@@
$\widetilde{p}_{g}=-\sum_{\alpha_{j}<0}\alpha_{j}.$
By the symmetry of the spectral numbers with respect to $\alpha\mapsto n-1-\alpha$, we can equivalently write
(2.4) $\widetilde{p}_{g}=\sum_{j}\lambda_{j}^{\prime},$
where the sum is now over rationals $10<\lambda_{j}^{\prime}<$ 1 such that $\exp(-2\pi i\lambda_{j}^{\prime})$ is an eigenvalue of $T_{s}$ acting on $\operatorname{Gr}_{F}^{0}H^{n}(\operatorname{Mil}_{f})$. This is to be compared with (2.2).
@@
(2.5) $\widetilde{p}_{g}=\sum_{\alpha_{j}^{\prime}<1}(1-\alpha_{j}^{\prime}).$
This convention appears in the works of M. Saito [34, 35], applied below. Depending on the context, one convention may be more adapted than the other, and we will use both.
2.4.
The Milnor number $\mu$ and the geometric genus $p_{g}$ depend only on the fiber $V=f^{-1}(0)$, but not on the chosen deformation [39, Theorem 2.9]. The spectral genus depends on the whole germ $f$. Nevertheless, the spectrum is constant in any deformation of isolated hypersurface singularities with constant $\mu$ [39, Theorem 2.8], and hence so is the spectral genus.
2.5.
@@
We also recall the Thom–Sebastiani property for the spectral numbers [41, Theorem 7.3]. If $h\colon(\mathbb{C}^{m+1},0)\to(\mathbb{C},0)$ defines another isolated hypersurface singularity, with spectral numbers $\beta_{j}^{\prime}$ taken in $(0,m+1)$, then the spectral numbers of $f(x_{0},\ldots,x_{n})+h(y_{0},\ldots,y_{m})$ are given by the sums
$\alpha_{i}^{\prime}+\beta_{j}^{\prime},\quad\text{for}\ i=1,\ldots,\mu_{f}\ \text{and}\ j=1,\ldots,\mu_{h},$
which hence belong to $(0,m+n+2)$.
2.6.
We next review some basic facts on the local and global theories of isolated singularities, in connection with degenerations of Hodge structures. Consider a connected complex manifold $\mathcal{X}$ of dimension $n+1\geq 2$, and a flat, projective morphism $g:\mathcal{X}\to\Delta$, which is a holomorphic submersion outside the origin. We will suppose that the central fiber has at most isolated singularities. If $x_{i}\in\mathcal{X}_{0}$ is such a singular point, the germ $(\mathcal{X},x_{i})\to(\Delta,0)$ is isomorphic to some $f_{i}:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$ and admits a Milnor fiber $\operatorname{Mil}_{f_{i}}$. In this setting, we denote by $\mu$ and $\widetilde{p}_{g}$ the sum of the Milnor numbers and spectral genera of the singularities $x_{i}$, respectively. Therefore, there is a decomposition
@@
for fixed $t\neq 0$, and for $q\neq n,n+1$ there is an isomorphism
(2.8) $H^{q}(\mathcal{X}_{0})\simeq H^{q}(\mathcal{X}_{t}).$
When the $H^{k}(\mathcal{X}_{0})$ are given the canonical mixed Hodge structures of Deligne, and the $H^{k}(\mathcal{X}_{t})$ are given the limit mixed Hodge structures of Schmid, then (2.7) is moreover an exact sequence of mixed Hodge structures. It is equivariant with respect to the semi-simple part of the monodromy, $T_{s}$. For details, we refer to Steenbrink [37, Section 3.3] and Navarro Aznar [26, Section 14]. Below, we will denote the limit mixed Hodge structure in degree $k$ simply by $(H^{k}_{\lim},F^{\bullet},W_{\bullet})$.
2.8.
@@
By an argument of Brieskorn [8, Section 1.1], such a compactification exists, and one can further suppose that it is given by a family of hypersurfaces in $\mathbb{P}^{n+1}$. This moreover shows that properties such as the positivity of $\frac{\mu}{(n+2)!}-\widetilde{p}_{g}$ can be studied equivalently for proper families or in the local setting. This will be used to reformulate our conjecture in Proposition 3.3 below.
@@ 3. Analytic torsion of $\mathcal{O}_{X}$
In this section, we discuss the asymptotic behavior of the holomorphic analytic torsion of the sheaf of holomorphic functions, for a degeneration of projective varieties, and relate it to our conjecture on the spectral genus.
@@
More generally, for a flat proper morphism of complex analytic spaces $\mathcal{X}\to S$ over a complex analytic manifold $S$, and a vector bundle $E$ on $\mathcal{X}$, there is a line bundle $\lambda(E)$ on $S$, whose fibers over $s\in S$ are given by $\lambda(E|_{\mathcal{X}_{s}})$, see [5, Section 4.1]. It is also referred to as the determinant of the cohomology. If $\mathcal{X}\to S$ is the analytification of an algebraic family, this construction is the analytification of the Knudsen–Mumford determinant [22]. It has a natural grading, which for the purposes of this article, together with various sign issues, we can ignore.
@@
Let $X$ be a compact Kähler manifold. The analytic torsion of a hermitian vector bundle $E$ on $X$ is defined as a weighted alternating product of determinants of Laplacians, namely
$\tau(X,E)=\exp\left(\sum(-1)^{q+1}q\zeta^{\prime}_{0,q}(0)\right)=\prod_{q}(\det\Delta_{\overline{\partial}}^{0,q})^{(-1)^{q}q}.$
Here, for $\operatorname{Re}(s)\gg 0$, $\zeta_{0,q}(s)$ is given by
@@
whose sum runs over positive eigenvalues $\mu_{j}$ of the Dolbeault-Laplacian $\Delta_{\overline{\partial}}^{0,q}$ acting on $A^{0,q}(E)$. It depends on both the Kähler metric on $X$ and the hermitian metric on $E$.
@@
Let $\mathcal{X}\to\Delta$ be a degeneration of projective varieties of dimension $n\geq 1$ as in §2.8 (1), with isolated singularities in the central fiber $\mathcal{X}_{0}$. Suppose that $\mathcal{X}$ is equipped with a Kähler metric, and that we are given a hermitian vector bundle $E$ on $\mathcal{X},$ of rank $e$. Let $\sigma$ be a holomorphic trivialization of the determinant of the cohomology $\lambda(E)$. By the main results of [46, 47], we have
(3.3) $\log\|\sigma\|_{Q}^{2}=\frac{(-1)^{n}}{(n+2)!}\mu\cdot e\cdot\log|t|^{2}+O(1),\quad\text{as}\quad t\to 0.$
@@
Suppose now that $\mathcal{X}\to\Delta$ is as in §3.4, but $\mathcal{X}$ is actually equipped with a Kähler metric, whose associated Kähler form is rational when restricted to smooth fibers. Endow the sheaf of holomorphic functions $\mathcal{O}_{\mathcal{X}}$ with the trivial hermitian metric, induced by the absolute value.
@@
$\log\|\sigma\|^{2}_{L^{2}}=(-1)^{n}\widetilde{p}_{g}\log|t|^{2}+(-1)^{n+1}\beta\log\log|t|^{-1}+O(1),$
where $\beta\geq 0$ is an integer determined by the limit mixed Hodge structure $H^{n}_{\lim}$, and given by
(3.4) $\beta=\sum_{r=1}^{n}r\dim F^{n}\operatorname{Gr}^{W}_{n+r}H^{n}_{\lim}.$
In particular, if $\mathcal{X}_{0}$ has canonical singularities, then $\log\|\sigma\|^{2}_{L^{2}}=O(1)$.
@@
First of all, the determinant of the cohomology of the sheaf of holomorphic functions is invariant under blowups in the special fiber, because the total space is smooth and hence has only rational singularities. We can hence suppose the central fiber of $\mathcal{X}\to\Delta$ has normal crossings. In this case, by [16, Theorem C] we find that
$\log\|\sigma\|_{L^{2}}^{2}=\left(\sum_{q=0}^{n}(-1)^{q}\alpha^{0,q}\right)\log|t|^{2}+\left(\sum_{q=0}^{n}(-1)^{q}\beta^{0,q}\right)\log\log|t|^{-1}+O(1),$
where $\alpha^{0,q}$ is defined as minus (the lower extension of) the logarithm of the semi-simple part of the monodromy acting on $\operatorname{Gr}_{F}^{0}H^{q}_{\lim}$, and
$\beta^{p,q}=\sum_{r=-k}^{k}r\dim\operatorname{Gr}_{F}^{p}\operatorname{Gr}^{W}_{k+r}H^{k}_{\lim},\quad\text{with}\quad k=p+q.$
One can infer directly from (2.4) and the exact sequence (2.7), that $\alpha^{0,n}=\widetilde{p}_{g}$, since the monodromy acts trivially on $H^{q}(\mathcal{X}_{0})$ for any $q$. Moreover, $\alpha^{0,q}=0$ if $q<n$. This latter fact follows from (2.7) and (2.8), paired again with the fact that $H^{q}(\mathcal{X}_{0})$ has trivial monodromy for any $q$. Similarly, since by (2.8) the mixed Hodge structure $H^{q}_{\lim}$ is pure for $q<n$, we see that $\beta^{0,q}=0$ for $q<n$. Finally, by [16, Lemma 4.3], $\beta^{0,n}=-\beta^{n,0}$, and we have $\beta^{n,0}=\beta$, because $\operatorname{Gr}_{F}^{n}=F^{n}$ and $F^{n}\operatorname{Gr}_{n+r}^{W}H^{n}_{\lim}=0$ if $r<0$. This concludes the proof.
For the second part of the proposition, we need to show that if $\mathcal{X}_{0}$ has canonical singularities, then $\widetilde{p}_{g}=\beta=0$. We already know that $p_{g}$, and hence $\widetilde{p}_{g}$, vanishes, cf. §2.2. The vanishing of $\beta$ is equivalent to a result of C.-L. Wang [42, Theorem 2.1 & Corollary 2.4], to the effect that the nilpotent operator $N$, associated to the monodromy on $H^{n}_{\lim}$, annihilates $F^{n}H^{n}_{\lim}$. Indeed, if $N$ annihilates $F^{n}H^{n}_{\lim}$, then the image of $F^{n}\operatorname{Gr}^{W}_{n+r}H^{n}_{\lim}$ under $N^{r}$ vanishes for $r\geq 1$, and we know that $N^{r}$ defines an isomorphism $\operatorname{Gr}^{W}_{n+r}H^{n}_{\lim}\to\operatorname{Gr}^{W}_{n-r}H^{n}_{\lim}$. Conversely, if $\beta=0$, then necessarily $F^{n}H^{n}_{\lim}=F^{n}W_{n}H^{n}_{\lim}$. But $N$ sends $F^{n}W_{n}H^{n}_{\lim}$ to $F^{n-1}W_{n-2}H^{n}_{\lim}$ and the latter vanishes, because $F^{n-1}\operatorname{Gr}^{W}_{r}H^{n}_{\lim}=0$ for every $r\leq n-2$, for type reasons.
We notice that the results in [42] require that $\mathcal{X}_{0}$ be Gorenstein and irreducible. The assumption that $\mathcal{X}_{0}$ is Gorenstein in op. cit. is automatic in our case, since $\mathcal{X}$ is smooth. As for the asumption that $\mathcal{X}_{0}$ is irreducible, it is not necessary in our setting. Indeed, since the morphism $\mathcal{X}\to\Delta$ has reduced fibers and $\mathcal{X}$ is smooth, the Stein factorization is of the form $\mathcal{X}\to\Delta^{\prime}\to\Delta$, where $\Delta^{\prime}$ is a disjoint union of discs and $\Delta^{\prime}\to\Delta$ is a trivial covering. Working over the components of $\Delta^{\prime}$ instead of $\Delta$, we reduce to the case that the fibers are connected. Since $\mathcal{X}_{0}$ has canonical singularities by assumption, it is in particular normal, so that connectedness entails irreducibility. ∎
@@
Let $\mathcal{X}\to\Delta$ be a degeneration as in §3.5. We conclude by (3.3) and Proposition 3.1 the following: